Thornton, Inc. needs to ship a 2,000 lbs of an industrial adhesive product to a customer. The dimensions of are height-35 inches, length = 40 inches, and width-48. It is a class 55. The LTL carrier offers a 55% discount. The cost per hundredweight (CWT) is $24.82. What is the total cost of shipment. Round to two decimal places. O $496.40 O $223.38 O$322.85 O Not enough information to answer
Answer #1

Weight = 2000 lbs

Cost per hundreweight = 24.82 $

Number of CWT = 2000/100 = 20

Total cost = 20*24.82 = 496.40

Discount = 55%

Total cost = 496.40*.40 = 223.38 $